# OpenFACADES: An Open Framework for Architectural Caption and Attribute Data Enrichment via Street View Imagery

### v1

@@ 4.2.4 Ablation experiments on data size
- - 33.09 28.97
5% 100% 42.46 41.82
10% 100% 42.02 41.42
25% 100% 41.51 41.00
50% 100% 40.47 40.20
75% 100% 40.81 40.32
100% 5% 34.98 35.72
100% 10% 35.86 36.18
100% 25% 38.25 38.20
100% 50% 39.14 38.83
100% 75% 39.58 39.33
100% 100% 40.28 39.87
In summary, for all structured variables, these results underscore a practical trade‐off: using only 5–10% of the OSM data (1,000-2,000 image–text pairs in this case) can achieve performance superior to the GPT‐4o baseline, while supplementing sufficient OSM data with minimal GPT-generated content can maximize model performance, thereby offering significant efficiency gains. Nonetheless, attaining state‐of‐the‐art descriptive captions likely requires more extensive textual annotations and larger model architectures to capture finer‐grained semantic and linguistic details. Striking a balance between captioning and labeling performance, we ultimately employ the model trained on the full dataset.
4.3 Comparative experiments
4.3.1 Computer vision models
As discussed in Section 2.2, CNNs and ViTs are widely used to infer building attributes. Here, we compare their performance with that of the fine‐tuned InternVL2.5‐2B VLM on four building characteristics, as summarized in Tables 7(a) and 7(b). To facilitate a fair comparison, we integrate GPT-generated data to supplement the missing OSM data in the training set for the CV models.
Table 7: Validation performance comparison between fine-tuned InternVL2.5-2B and CV models.
Attribute Model Accuracy Precision Recall F1 Acc@2
Building type DenseNet 0.54 0.54 0.54 0.53 0.68
VGG 0.47 0.47 0.47 0.47 0.64
ResNet 0.52 0.54 0.52 0.52 0.67
ResNet18 0.51 0.51 0.51 0.51 0.67
ResNet101 0.54 0.55 0.54 0.53 0.67
ViT16 0.54 0.56 0.54 0.54 0.69
ViT32 0.52 0.52 0.52 0.51 0.68
InternVL2.5-2B 0.61 0.64 0.61 0.60 0.76
Surface material DenseNet 0.65 0.67 0.65 0.64 0.81
VGG 0.57 0.61 0.57 0.57 0.74
ResNet 0.65 0.67 0.65 0.64 0.79
ResNet18 0.61 0.64 0.61 0.61 0.79
ResNet101 0.66 0.69 0.66 0.65 0.81
ViT16 0.65 0.68 0.65 0.65 0.79
ViT32 0.63 0.67 0.63 0.64 0.78
InternVL2.5-2B 0.69 0.74 0.69 0.68 0.82
Attribute Model R2 ($\uparrow$) MAE ($\downarrow$) MAPE ($\downarrow$) RMSE ($\downarrow$)
Number of floors DenseNet 0.77 2.35 0.40 4.27
VGG 0.67 3.15 0.42 5.46
ResNet50 0.78 2.44 0.40 4.54
ResNet18 0.75 2.77 0.40 4.72
ResNet101 0.77 2.52 0.41 4.60
ViT16 0.77 2.44 0.38 4.52
ViT32 0.75 2.58 0.45 4.72
InternVL2.5-2B 0.77 2.22 0.35 4.53
Building age DenseNet 0.72 30.80 1.03 50.26
VGG 0.56 41.65 1.35 63.13
ResNet50 0.72 32.34 1.10 50.14
ResNet18 0.68 34.66 1.11 53.50
ResNet101 0.72 31.25 1.01 50.04
ViT16 0.71 31.82 1.20 50.82
ViT32 0.68 34.45 1.25 53.51
InternVL2.5-2B 0.70 28.64 0.65 51.67
In general, the fine-tuned VLM (InternVL2.5‐2B) achieves the best performance, particularly in tasks such as building type and surface material predictions. While CV models slightly exceed the performance of VLMs in predicting the number of floors and building age when trained on ChatGPT-generated data, these models require domain-specific tuning and separate architectures for each attribute. VLMs, however, offer a unified and highly adaptable approach, achieving comparable or superior performance across multi-attribute prediction tasks. This underscores the advantage of leveraging pretrained semantic reasoning and contextual understanding inherent in VLMs, which generalize well across diverse tasks.
Additionally, VLMs provide the capability to generate textual descriptions of architectural and contextual details, offering richer insights into building attributes and enabling further qualitative or descriptive analyses. This combination of robust predictive performance and expanded functionality makes VLMs a compelling alternative to traditional approaches, particularly in applications requiring adaptability and scalability across complex urban environments.
4.3.2 Generalizability
Generalizing CV models to unseen cities remains a significant challenge due to the diverse and unique architectural features across cities (Sun et al., 2022b). VLMs, pretrained on extensive, high‐quality image‐text datasets, demonstrate promising potential to overcome these limitations by leveraging their pre‐acquired semantic reasoning and contextual understanding capabilities. To investigate this potential, we conducted an experiment on building imagery collected from Brussels, comparing the performance of established CNN and ViT architectures against our fine‐tuned VLM. For this evaluation, we curated a dataset of 3,728 labeled building images by integrating OSM attributes with buildings detected from Mapillary SVI. The dataset is composed of 3,383 images for building type, 186 for surface material, 1,245 for the number of floors, and 99 for building age.
Table 8(a) and 8(b) indicate that the VLM model demonstrates superior generalizability compared to commonly used CV models, which is particularly evident in the tasks of building type and age prediction. The enhanced performance can be attributed to the pre-trained VLM’s ability to leverage semantic reasoning and contextual understanding from its large-scale image-text pretrain data, enabling it to adapt to diverse architectural features. While performance is similar for number of floors predictions, the lower accuracy of ResNet101 and ViT16 on other attributes highlights their reliance on localized visual features, limiting generalization. For surface material classification, VLM achieves high precision, but overall performance is affected by label inconsistencies in the dataset. For instance, buildings labeled as “brick” in OSM are often visually identified as “plaster” during manual inspection, a discrepancy linked to local labeling conventions in Brussels. This issue reflects broader challenges with data quality and the discrepancies in labeling schemes across cities. Developing an unified or context-aware annotation system would be valuable for addressing such inconsistencies.
Attribute Model Accuracy F1 Precision Recall
Building type ResNet101 0.37 0.46 0.64 0.37
ViT16 0.26 0.35 0.68 0.26
InternVL2.5-2B 0.58 0.63 0.74 0.58
Surface material ResNet101 0.19 0.31 0.93 0.19
ViT16 0.20 0.32 0.84 0.20
InternVL2.5-2B 0.31 0.44 0.93 0.31
Attribute Model RMSE ($\uparrow$) MAE ($\downarrow$) MAPE ($\downarrow$) R2 ($\uparrow$)
Number of floors ResNet101 1.91 1.12 0.32 0.57
ViT16 1.78 1.06 0.29 0.63
InternVL2.5-2B 1.88 0.94 0.27 0.59
Building age ResNet101 50.72 39.05 1.81 0.15
ViT16 49.36 36.81 1.93 0.22
InternVL2.5-2B 46.03 31.66 1.50 0.44
4.3.3 Sensitivity and robustness
As described in Section 3.3, we evaluate the robustness of the VLM against image corruptions by testing it on a perturbed dataset derived from the test set. Figure 12 illustrates the model’s performance under varying severity levels of occlusion, motion blur, Gaussian noise, and brightness distortions. In general, the model demonstrates resilience, with performance dropping by less than 10% under most mild and moderate image corruptions. In particular, the model remains significantly stable in handling lighting variations and occlusion, both of which are common challenges in crowdsourced image datasets. However, the model experiences a significant performance drop when confronted with moderate to severe noise and blurriness. In particular, the model is most affected by motion blur, where the error rate for the number of floors prediction increases from 0.25 (clean error) to 0.93, and the error rate for building age prediction rises from 0.29 (clean error) to 1.48. These findings emphasize the necessity of image preprocessing techniques to filter out degraded images during the image selection stage.
Furthermore, the Relative $m$CE is computed for all models using ResNet50 as the baseline. Table 9 presents the relative error rates across different building attribute prediction tasks, indicating each model’s stability compared to the baseline. In general, different models demonstrate various capability in handing corruption data. Multi-attribute prediction VLM (InternVL2.5-2B) demonstrates superior stability compared to single-attribute CV models in most cases, especially when distinguishing building type and surface material and when encountering occlusion and brightness variations. CNN models’ stability performs comparably to more advanced models in the tasks of number of floors prediction, while ViT model performs superior in handling data with Gaussian noise. This outcome implies that additional domain-specific constraints or specialized training strategies might be required to enhance performance on crowdsourced image data.
Table 9: Relative prediction error under different corruptions and perturbations for different objective attributes.
Attribute Model Occlusion Motion Noise Brightness Relative mCE
Building type ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 0.95 1.14 0.90 1.10 1.02
ViT16 1.09 1.22 0.59 1.60 1.13
InternVL2.5-2B 0.51 0.65 0.80 0.64 0.65
Surface material ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 1.00 0.84 0.94 0.78 0.89
ViT16 1.61 1.01 0.79 1.64 1.26
InternVL2.5-2B 0.45 0.59 0.67 0.51 0.56
Number of floors ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 1.05 0.91 0.82 0.89 0.92
ViT16 0.95 0.82 0.27 2.05 1.02
InternVL2.5-2B 1.07 1.44 0.86 0.85 1.05
Building age ResNet50 1.00 1.00 1.00 1.00 1.00
ResNet101 0.58 0.85 0.95 0.81 0.80
ViT16 0.40 0.62 0.70 1.99 0.93
InternVL2.5-2B 0.15 0.72 1.44 0.56 0.72
In conclusion, building on insights from prior CV techniques, VLMs not only demonstrate robust and generalizable features for tackling diverse tasks based on crowdsourced data, but they also represent a promising framework for large-scale or cross-regional implementations that demand multi-feature prediction and flexible adaptation. Moreover, incorporating more domain-specific and diverse, high-quality data can further enhance the framework’s performance and facilitate broader adoption in real-world scenarios.
@@ 4.4 Image labeling and captioning
Detected buildings across seven global cities, introduced in Section 4.1, are subsequently processed by the fine‐tuned VLM to generate objective attributes and captions. Overall, data for half a million buildings are enriched using 1.2 million images, each linked to its geographical location. For buildings with multiple observations, the most frequently assigned categories are retained. Figure 13 compares the availability of building properties before and after enrichment in Washington, D.C. The proposed approach effectively enhances building‐level information, particularly for surface material and building age, wherever street‐level imagery is available. Additionally, Table 10(d) presents the distribution of class labels for each attribute across the 1.2 million‐building dataset.
@@ 5.1 Application of building image dataset
Despite the centrality of objective building attributes in urban analytics, their scarcity still persists across cities (Biljecki et al., 2023). Our open framework OpenFACADES, addresses this gap by utilizing SVI — an urban sensing modality that captures pedestrian-scale visual information — together with building data to develop an open-sourced MLLM framework for unified attribute extraction and semantic description. The methodology begins by integrating crowdsourced SVI metadata with geometrical building data using isovist analysis to identify relevant images. Buildings are then detected based on their angles of view within image space, followed by an automated process of reprojecting and filtering them into individual building images. Lastly, a subset of this dataset is used to construct an image-text dataset designed for three tasks for VLM fine-tuning: single-word Q&A, multi-attribute Q&A, and captioning. Our experiments indicate that the fine-tuned VLM demonstrates strong performance in multi-attribute prediction, surpassing CV models and outperforming zero-shot ChatGPT-4o baselines. Deploying the VLM at scale, we annotate and release data of half a million buildings with both objective attributes and textual descriptions, derived from 1.2 million images across seven global cities, contributing to a scalable and automated approach for building property enrichment.
Our study directly features three main contributions to building research. First, our methodology detects holistic building facades and reprojects them into undistorted individual images, ensuring comprehensive visual coverage while reducing the uncertainty inherent in panoramic imagery. This pipeline can be integrated with existing methods to detect buildings from diverse viewing angles and associate them with geolocation, enabling nuanced and holistic observation for exterior modeling (Zhang et al., 2021), facade material segmentation (Tarkhan et al., 2025), and window-to-wall ratio calculation (De Simone et al., 2024). Second, this work introduces an inclusive and efficient pipeline to utilize both crowdsourced data and open-sourced LLMs for street-level research. This pipeline not only overcomes the challenge of relying on proprietary datasets, but also circumvents the high costs and limited adaptability associated with proprietary LLM APIs, making advanced analytical techniques more accessible and reproducible to the research community. Future studies might adjust the pipeline to customized tasks to incorporate fine-grained visual information with tailored building data based on their objectives, such as building conditions (Zou and Wang, 2021), human perceptual indicators (Liang et al., 2024) and seismic structural types (Pelizari et al., 2021).
@@
  • Explore the capabilities of VLMs through zero‐shot settings, varying scales of training data, and applications across different cities and attribute categories. Our findings show that fine‐tuning is essential to enhance performance in current VLMs, especially given the inherent quality issues in crowdsourced building data. Augmenting computer‐generated labels with ground‐truth annotations significantly improves model performance. Furthermore, our experiments demonstrate that efficient training can be achieved by using only 5% of the available crowdsourced data (approximately 1,000 image–text pairs), or by supplementing sufficient ground truth building data with minimal GPT-generated content, while still yielding robust performance gains of multi-attribute prediction.
@@
  • Apply the method at scale by generating labels and captions for half a million buildings in eight cities, laying a foundation for future urban analyses. For instance, integrating these labeled data with geospatial information can add new dimensions to urban functional zone classification (Zhang et al., 2023), including potential insights into 3D functional zoning (Lin et al., 2024). The unified model also infers multi‐dimensional building properties relevant for applications such as modeling building electricity consumption (Rosenfelder et al., 2021), estimating material stocks (Raghu et al., 2023), and assessing structural risk (Wang et al., 2021). Additionally, captions offer an extra layer of information about building facades, enabling the identification of mixed‐use buildings or stylistic variations. This linguistic data holds promise for exploring urban identity, supporting text‐image‐based generative design, and serving as an additional feature layer in multimodal model training.
  • Demonstrate enhanced generalizability and robustness by comparing VLMs with CV models. Owing to larger parameter counts and extensive pretraining on general‐purpose data, VLMs excel in predicting objective building attributes for both unseen data from culturally different region and images with variable conditions. This result provides a convincing avenue for future cross‐city analyses and crowdsourced data research, where broad adaptation and resilience to heterogeneous imagery remain critical challenges.
5.2 Limitations and future works
@@
Despite the advancements presented in this study, limitations remain. First, while this study incorporates captioning data for fine-tuning VLMs, these captions are generated using commercial state-of-the-art LLMs rather than human-labeled ground truth, leaving their accuracy and reliability unverified. A systematic human evaluation would be valuable for future research to assess captioning quality, consistency, and semantic accuracy. Additionally, leveraging open-access models offers a more sustainable approach for scalable dataset expansion in future studies. Knowledge distillation — where a smaller model learns from a larger teacher model — presents a promising solution for self-supervised learning, enabling broader generalization across diverse urban settings and more efficient adaptation to building-related tasks.
@@
Second, while the fine-tuned model exhibits strong generalizability across cities, the quality of crowdsourced data remains a crucial factor (Biljecki et al., 2023; Hou and Biljecki, 2022). Although this study incorporates multiple strategies to mitigate data quality issues — such as isovist-based vantage point selection, building data harmonization, and feature-based image filtering — challenges persist. These include incorrect or incomplete building labels, inconsistent geometry information, non-standardized image format, and misaligned image coordinates, all of which contribute to different sources of uncertainty. Future work should focus on enhancing dataset reliability through improved data filtering mechanisms. Automated repetition detection, heuristic rule-based filtering, and uncertainty-aware sampling could refine image selection and minimize inconsistencies in building attribute annotations (Chen et al., 2024b).
@@ 6 Conclusion
This comprehensive study advances spatial data infrastructures and urban data science by introducing a novel framework, OpenFACADES, which leverages volunteered geographic information to enrich building profiles on a global scale using street-level imagery and multimodal large language models. We harvest multimodal crowdsourced data and apply isovist analysis, object detection, and a tailored reprojection method to geolocate and acquire holistic building images, thereby establishing a comprehensive global building image dataset. A selection of this open dataset is then utilized for fine-tuning large vision-language models (VLMs), enabling large-scale enrichment of building profiles through multi-attribute prediction (e.g., material, function) and open-vocabulary captioning.
Our findings demonstrate that VLMs outperform conventional CNN-based models and zero-shot GPT-4o baselines in predicting building attributes while generating linguistically grounded descriptions. This methodological advancement has enabled the creation of a large-scale dataset covering half a million buildings across seven global cities, addressing critical gaps in building data availability for urban analytics. By bridging the limitations of existing datasets, this framework provides a scalable solution for capturing multi-dimensional fine-grained architectural details and urban morphological characteristics. The enriched dataset further facilitates a more nuanced and expansive exploration of urban environments, with potential applications in energy modeling, risk assessment, and sustainable development.

### v2

@@ 5.4 Ablation experiments
- - 36.35 22.54 15.11 10.16 28.32 27.11
5% 100% 57.17 42.09 33.03 26.31 44.97 43.53
10% 100% 57.50 42.40 33.33 26.59 45.29 43.74
25% 100% 57.25 42.17 33.12 26.39 45.07 43.63
50% 100% 56.71 41.73 32.75 26.09 44.62 43.35
75% 100% 56.97 41.92 32.89 26.18 44.88 43.39
100% 5% 53.48 37.78 28.57 21.87 41.12 39.25
100% 10% 54.36 38.62 29.37 22.64 41.67 40.02
100% 25% 54.06 38.75 29.71 23.08 41.78 40.65
100% 50% 55.24 40.13 31.13 24.50 43.02 41.92
100% 75% 55.92 41.26 32.08 25.59 44.09 42.84
100% 100% 56.84 41.82 32.81 26.13 44.81 43.40
In summary, our ablations show that (1) full fine-tuning of open-source MLLMs delivers stronger domain adaptation than adapter-based methods, with fewer epochs and smaller learning rates achieving comparable knowledge gains; (2) small-scale models strike the best efficiency–performance balance, while larger models maintain an edge in free-form captioning; and (3) a modest, well-curated dataset captures most of the benefits of large-scale pretraining, although over-emphasis on structured labels can modestly reduce generative richness. Together, these findings underscore the importance of harmonizing fine-tuning scope, model capacity, data quality, and training objectives when extending vision–language models to building-specific tasks.
6 Discussion
@@ 6.1 Image labeling and captioning
Detected buildings across seven global cities, introduced in Section 5.1, are subsequently processed by the fine‐tuned VLM to generate objective attributes and captions. Overall, data for half a million buildings are enriched using 1.2 million images, each linked to its geographical location. For buildings with multiple observations, the most frequently assigned categories are retained. Figure 14 compares the availability of building properties before and after enrichment in Washington, D.C. The proposed approach effectively enhances building‐level information, particularly for surface material and building age. The supplementary material presents the distribution of class labels for each attribute across the 1.2 million‐building dataset.
@@ 6.2 Application of building image dataset
Despite the centrality of objective building attributes in urban analytics, their scarcity still persists across cities (Biljecki et al., 2023). Our open-source framework OpenFACADES addresses this gap by combining SVI, which captures pedestrian-scale visuals, with building data to train an MLLM for unified attribute extraction and semantic description. The methodology begins by integrating crowdsourced SVI metadata with geometrical building data using isovist analysis to identify relevant images. Buildings are then detected based on their angles of view within image space, followed by an automated process of reprojecting and filtering them into individual building images. Lastly, a subset of this dataset is used to construct an image-text dataset designed for three tasks for VLM fine-tuning: single-attribute Q&A, multi-attribute Q&A, and captioning. Our experiments indicate that the fine-tuned VLM demonstrates strong performance in multi-attribute prediction, surpassing CV models and outperforming zero-shot ChatGPT-4o baselines. Deploying the VLM at scale, we annotate and release data of half a million buildings with both objective attributes and textual descriptions, derived from 1.2 million images across seven global cities, contributing to a scalable and automated approach for building property enrichment.
Our study directly features three main contributions to building research. First, our methodology detects holistic building facades and reprojects them into undistorted individual images, ensuring comprehensive visual coverage while reducing the uncertainty inherent in panoramic imagery. This pipeline can be integrated with existing methods to detect buildings from diverse viewing angles and associate them with geolocation, enabling nuanced and holistic observation for exterior modeling (Zhang et al., 2021), facade material segmentation (Tarkhan et al., 2025), and window-to-wall ratio calculation (De Simone et al., 2024). Second, this work introduces an inclusive and efficient pipeline to utilize both crowdsourced data and open-sourced LLMs for street-level research. This pipeline not only overcomes the challenge of relying on proprietary datasets, but also circumvents the high costs and limited adaptability associated with proprietary LLM APIs, making advanced analytical techniques more accessible and reproducible to the research community. Future studies might adjust the pipeline to customized tasks to incorporate fine-grained visual information with tailored building data based on their objectives, such as building conditions (Zou and Wang, 2021; Zhang et al., 2024b), human perceptual indicators (Liang et al., 2024) and seismic structural types (Pelizari et al., 2021).
@@
  • Demonstrate that full fine-tuning of a open-source VLM backbone yields state-of-the-art multi-attribute extraction. It surpass zero-shot ChatGPT and matching or exceeding specialized CNN/ViT baselines on building type, material, floor count and age, simultaneously generating coherent captions within one unified framework. The model delivers consistent performance across categories and cities.
  • Reveal generalizability and robustness. In cross-city evaluations and synthetic corruption tests, our fine-tuned VLM surpasses CV models by maintaining high accuracy on unseen urban contexts and showing resilience to occlusion, blur, and lighting distortions, underscoring its suitability for heterogeneous, crowdsourced SVI.
  • Investigate efficient training paradigms. Through systematic ablations, we show that (i) full fine-tuning of open-source VLMs with a low learning rate over a few epochs, (ii) the use of a small-scale backbone, and (iii) a well-curated, balanced dataset together capture most benefits of large-scale pretraining by minimizing data, compute, and architectural complexity while preserving accurate multi-attribute prediction and high-quality captioning.
@@
Finally, we apply the pipeline at scale by generating labels and captions for half a million buildings in seven cities, laying a foundation for future urban analyses. For instance, integrating these labeled data with geospatial information can add new dimensions to urban functional zone classification (Zhang et al., 2023), including potential insights into 3D functional zoning (Lin et al., 2024). The unified model also infers multi‐dimensional building properties relevant for applications such as modeling building electricity consumption (Rosenfelder et al., 2021), estimating material stocks (Raghu et al., 2023), and assessing structural risk (Wang et al., 2021). Additionally, captions offer an extra layer of information about building facades, enabling the identification of mixed‐use buildings or stylistic variations. This linguistic data holds promise for exploring urban identity, supporting text‐image‐based generative design, and serving as an additional feature layer in multimodal model training.
6.3 Limitations and future work
@@
Despite the advancements presented in this study, limitations remain. First, while this study incorporates captioning data for fine-tuning VLMs, these captions are generated using commercial state-of-the-art LLMs rather than human-labeled ground truth, leaving their accuracy and reliability unverified. A systematic human evaluation would be valuable for future research to assess captioning quality, consistency, and semantic accuracy. Additionally, leveraging open-access models offers a more sustainable approach for scalable dataset expansion in future studies. Knowledge distillation, in which a smaller model learns from a larger teacher, offers a promising self-supervised approach to improve generalization across urban contexts.
@@
Second, while the fine-tuned model exhibits strong generalizability across cities, the quality of crowdsourced data remains a crucial factor (Biljecki et al., 2023; Hou and Biljecki, 2022). Although multiple strategies were employed in this study to mitigate data quality issues, several challenges persist. These include incorrect or incomplete building labels, inconsistent geometric information, non-standardized image formats, and misaligned image coordinates, each contributing to various uncertainties. Additionally, as no manual image selection was performed in this study, potential biases in data collection remain unaddressed. Future work should focus on enhancing dataset reliability through improved data filtering mechanisms. Automated repetition detection, heuristic rule-based filtering, and uncertainty-aware sampling could refine image selection and minimize inconsistencies in building attribute annotations (Chen et al., 2024b).
@@ 7 Conclusion
This comprehensive study advances spatial data infrastructures and urban data science by introducing a novel framework, OpenFACADES, which leverages volunteered geographic information to enrich building profiles on a global scale using street-level imagery and multimodal large language models. We harvest multimodal crowdsourced data and apply isovist analysis, object detection, and a tailored reprojection method to geolocate and acquire holistic building images, thereby establishing a comprehensive global building image dataset. A selection of this open dataset is then utilized for fine-tuning VLMs, enabling large-scale enrichment of building profiles through multi-attribute prediction and open-vocabulary captioning. This framework provides a scalable solution for capturing multi-dimensional fine-grained architectural details and urban morphological characteristics.
Our findings also demonstrate that VLMs generally outperform conventional CNN-based models and zero-shot GPT-4o baselines in predicting building attributes while generating linguistically grounded descriptions. This methodological advancement has enabled the creation of a large-scale dataset covering half a million buildings across seven global cities. The enriched dataset further facilitates a more nuanced and expansive exploration of urban environments, with potential applications in energy modeling, risk assessment, and sustainable development.